OpenStudy (anonymous):

FINAL REVIEW HELP: multiply the scientific notation: (6 x 10^-7)(5.3 x 10^-7)

OpenStudy (phi):

mult the numbers. you can re-arrange them as \[ 6 \cdot 5.3 \cdot 10^{-7} \cdot 10^{-7} \]

OpenStudy (phi):

If the final answer is out side the range 1 to 9.999, move the decimal point and adjust the exponent

OpenStudy (anonymous):

is the answer 3.18 x 10^49

OpenStudy (phi):

you add exponents. remember 10^7 (for example) means 10*10*10*10*10*10*10 10 times itself 7 times and 10^7 * 10^7= 10*10*10*10*10*10*10 * 10*10*10*10*10*10*10 you can count the 10's but it is easier to add 7+7=14 It still works when the 7 is -7 remember \[ 10^{-7} \cdot 10^{-7} = \frac{1}{10^7}\cdot \frac{1}{10^7} = \frac{1}{10^7\cdot 10^{7}} \]

OpenStudy (anonymous):

oh, so is the answer 3.18 x 10^-13

OpenStudy (phi):

yes. \[6 \cdot 5.3 \cdot 10^{-7} \cdot 10^{-7} = 31.8 \cdot 10^{-14} \] 31.8 is ≥ 10, so divide it by 10 and multiply by 10: \[ \frac{31.8}{10} \cdot 10^{-14}\cdot 10^1 \] which simplifies to \[ 3.18 \cdot 10^{-13} \]